OpenStudy (anonymous):

Brendon Butterworth brought a present for his mother. He spent $28 plus 8% sales tax. Find the total cost of the present. Can Somebody show me the process of figuring out questions like this one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

method one find 8% of 28 and add to 28 you get \[.08\times 28=2.24\] and \[28+2.24=30.24\]

OpenStudy (anonymous):

method 2 multiply 28 by 1.08 to get \[28\times 1.08=30.24\] second method best if you have a calculator

OpenStudy (anonymous):

Thank you that really helped!